OpenStudy (anonymous):

Suppose y= (x/sqrt(a^2-1) ) - (2/sqrt(a^2-1)) * tan^-1(sinx/a+sqrt(a^2-1)+cosx) Show y`=1/a+cosx Please explain how you got it also! Suppose y= (x/sqrt(a^2-1) ) - (2/sqrt(a^2-1)) * tan^-1(sinx/a+sqrt(a^2-1)+cosx) Show y`=1/a+cosx Please explain how you got it also! @Mathematics

OpenStudy (anonymous):

\[y = \left( \frac{ x }{ \sqrt{a ^{2} - 1} } \right) - \left( \frac{ 2 }{ \sqrt{a ^{2} -1}} \right) \times \tan^{-1} \left( \frac{ \sin x }{ a } + \sqrt{a ^{2} -1} + \cos x\right)\] is this ur equation?

OpenStudy (anonymous):

yes